Part I Use the following information to answer the questions below. ObservationPriceQuantity    A $4.00 16   ...

Part I


Use the following information to answer the questions below.


ObservationPriceQuantity
   A $4.00 16
   B $6.00 10

1. Calculate a price elasticity of demand.  You must show all your work to earn credit.

2. Given the elasticity of demand, a 10% increase in price will cause quantity demanded to fall by what percentage?  Explain your answer.

3. Is this demand elastic or inelastic?  Explain your answer.

Part II

Walmart advertises that it has rolled back prices. If Walmart is rolling back prices to raise revenues, should it roll back prices on products that have a price elasticity of demand that is elastic or inelastic.  Explain your answer.
Answer #1

Part 1

1). Price Elasticity of Demand is the percentage change in quantity demanded due to a percentage change in price.

Elasticity = % Change in quantity demanded / % Change in price

% change in quantity demanded = (16 - 10) / 16 * 100 = 37.5%

% change in price = (6 - 4) / 4 * 100 = 50%

Elasticity = 37.5/ 50 = 0.75

Price elasticity of demand is 0.75 that means change of 1% in price will bring 0.75% change in quanitty demanded.

2). We have calculated price elasticity of demand in the above question, now we will use price elasticity formula to find change in quantity demanded ,

Elasticity = % Change in quantity demanded / % Change in price

0.75 = % change in quantity demanded / 10

% change in quantity demanded = 7.5%

that means when there is a 10% rise in price, demand will reduce by 7.5%.

3).

Price Elasticity Value
Elastic greater than 1
Unitary Equal to 1
Inelastic Less than 1

When elasticity is greater than 1, then demand is elastic because the change in demand is more than the change in price.

When elasticity is equal to 1, then demand is unitary elastic, that means the change in demand is equivalent to the change in price.

When elasticity is less than 1, then demand is inelastic, that means the change in demand is less than the change in price.

In the give question, our price elasticity is 0.75 which is less than 1 hence our demand is inelastic. Here, a change of 1% in the price, will bring a 0.75% change in the quantity demanded.

Part 2

We know that total revenue is equivalent to price multiplied by quantity. When Walmart is planning to roll back prices i.e to reduce the price, then it will raise revenue only when the % fall in prices creates a larger % increase in demand, so that the total revenue which is the product of both, would increase. So, % change in demand would be more than % change in price only when demand is elastic. Hence, if Walmart wants to increase the revenue, they should roll back the price of the goods whose demand is elastic.

We can understand it better through the below example,

In the above table, we have taken a 10% fall in price and a 5% increase in quantity when demand is inelastic (0.5) and 20% increase in quantity when demand is elastic (2). So, we can see that when our demand is elastic our total revenue has increased to 10,800 whereas it decreased to 9450 when demand was inelastic.